### v2

# Redshift leverage for the search of GRB neutrinos affected by quantum properties of spacetime

Abstract
Some previous studies based on IceCube neutrinos had found intriguing preliminary evidence that some of them might be GRB neutrinos with travel times affected by quantum properties of spacetime delaying them proportionally to their energy, an effect often labeled as “quantum-spacetime-induced in-vacuo dispersion”. Those previous studies looked for candidate GRB neutrinos in a fixed (neutrino-energy-independent) time window after the GRB onset and relied rather crucially on crude estimates of the redshift of GRBs whose redshift has not been measured. We here introduce a complementary approach to the search of quantum-spacetime-affected GRB neutrinos which restricts the analysis to GRBs of sharply known redshift, and, in a way that we argue is synergistic with having sharp information on redshift, adopts a neutrino-energy-dependent time window. We find that knowing the redshift of the GRBs strengthens the analysis enough to compensate for the fact that of course the restriction to GRBs of known redshift reduces the number of candidate GRB neutrinos. And rather remarkably our estimate of the magnitude of the in-vacuo-dispersion effects is fully consistent with what had been found using the previous approach. Our findings are still inconclusive, since their significance is quantified by a $p$-value of little less than $0.01$, but provide motivation for monitoring the accrual of neutrino observations by IceCube and KM3NeT as well as for further refinements of the strategy of analysis here proposed.
I Introduction and new selection criteria
The current generation of neutrino telescopes was devised to mark the start of neutrino astrophysics, and indeed IceCube firmly established the observation of cosmological neutrinos [1], but so far not much astrophysics has been done with neutrinos: there is only evidence of a neutrino signal from the active galaxy NGC 1068 [2] and from the flaring blazar TXS 0506+056 (see, e.g., Refs. [3, 4]). With respect to pre-IceCube expectations what is clearly missing are observations of neutrinos from GRBs (gamma-ray bursts). The prediction of a neutrino emission associated with GRBs is generic within the most widely accepted astrophysical models [5] and pre-IceCube studies estimated about a handful of GRB-neutrino observations per year of operation [6, 7, 8, 9]; yet, as IceCube nears 15 years of operation, still not a single GRB neutrino has been reported. The simplest (and most likely) explanation of this huge underperformance of IceCube is that our GRB models must be revised in such a way to produce a much lower neutrino flux, but it is intriguing that in principle our failure to observe GRB neutrinos could also be attributed to some plausible quantum properties of spacetime which had already been investigated for independent reasons in the quantum-gravity literature: IceCube searches might have failed to find GRB neutrinos because they assume that a GRB neutrino should be detected in very close temporal coincidence with the associated gamma rays, but a sizable mismatch between the neutrino detection time and the time of detection of the electromagnetic GRB signal would be expected in presence of “in-vacuo dispersion”, a much-studied effect such that quantum properties of spacetime slow down [^1] While the theoretical prejudice of most quantum-spacetime researchers favors effects that slow down particles, there is no conclusive theoretical argument ruling out that quantum-spacetime effects might instead speed up particles. We offer some data-based observations relevant for this issue in parts of Sec. V. particles proportionally to their energies (see, e.g., Refs. [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and references therein).
We here focus exclusively on the most studied scenario for in-vacuo dispersion which for our purposes is conveniently characterized in terms of the following relationship [21, 22, 23, 24, 20]:
$\Delta t=\eta D(z)\frac{E}{M_{P}}\,,$ (1)
where $\Delta t$ is the contribution to the travel time of the neutrino due to quantum-gravity-induced in-vacuo dispersion, $E$ is the neutrino energy, $M_{P}$ denotes the Planck scale ($\sim 10^{16}$ TeV), and $D(z)$ is a function of the redshift $z$ of the source,
[…]
($\Omega_{\Lambda}$, $H_{0}$ and $\Omega_{m}$ denote, as usual, respectively the cosmological constant, the Hubble parameter and the matter fraction, for which we take the values given in Ref. [25]).
We restrict to the scenario with purely systematic effects, in which $\eta$ is simply a parameter to be determined experimentally. When a larger collection of astrophysical neutrinos becomes available (and therefore analyses such as ours can attempt multi-parameter investigations), it will be possible to consider the more general case of a mixture of systematic and fuzzy effects [18, 19, 26], in which $\eta$ is described as a random variable (taking different values for different neutrinos) with mean $\eta_{0}$ and dispersion $\delta\eta$. We also postpone the extension of our phenomenological model to allow for the possibility that there might be intrinsic at-the-source effects causing a sizable offset between the typical neutrino emission time and the GRB onset. A first step for modeling the impact of such intrinsic effects on in-vacuo dispersion analyses would be to add to Eq.(1) an extra term of the form ${\bar{\tau}}(1+z)$, where $\bar{\tau}$ is the average neutrino emission time offset with respect to the GRB onset and the factor $(1+z)$ accounts for ordinary time dilation, but investigations of such a refined model need more astrophysical neutrinos than presently available.
The effects of Eq. (1) are totally unnoticeable on terrestrial scales, but for the observation of distant astrophysical sources the factor $D(z)$ can be large enough to compensate for the Planck-scale suppression [10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
Indeed, in-vacuo dispersion for photons has been investigated very intensely over the last 15 years (see, e.g., Refs. [18, 19] and references therein), with some studies even reaching sensitivity of the order of the Planck scale (see, e.g., Ref. [27]). It is difficult to say exactly which values of $\eta$ have been excluded for photons, since the magnitude of in-vacuo-dispersion effects could be comparable to known (but poorly understood) intrinsic spectral lags and a delayed emission of the relevant photons cannot be ruled out (some of the most energetic photons from a GRB could be emitted within the long afterglow phase). Still, a balanced perspective is that, for photons, we are close to Planck-scale sensitivity [18, 19].
The investigation of in-vacuo dispersion for neutrinos has been lagging far behind: the current bound on in-vacuo dispersion for neutrinos is still based on the historic observation of neutrinos from the SN1987a supernova, and only amounts to $\eta\lesssim 10^{9}$ [28]. This is very unsatisfactory, especially in light of the fact that the quantum-spacetime literature provides motivation for studying in-vacuo dispersion for photons and neutrinos separately. In particular, quantum-spacetime models based on Planck-scale discretization predict in-vacuo dispersion for neutrinos but not for photons [29, 30], and this is consistent with results from models based on causal sets [31]. Moreover, within a popular effective-theory approach to quantum-spacetime properties one describes in-vacuo dispersion for neutrinos in a way that is completely independent from that applicable to photons [32]. Given the typical energies of the astrophysical neutrinos observed by current telescopes, even just a single observation of a GRB neutrino would allow us to constrain in-vacuo dispersion for neutrinos rather tightly. However, as mentioned at the beginning, no such observation has been reported yet by studies assuming no in-vacuo dispersion.
As observed in Refs. [21, 22, 23, 24, 20, 33, 34] the search of GRB neutrinos affected by in-vacuo dispersion must necessarily rely on a statistical approach: we could never be sure that a certain neutrino is a GRB neutrino, but we might at some point observe enough GRB-neutrino candidates (neutrinos that could come from a GRB, assuming in-vacuo dispersion) to be sure that at least some of them are actually GRB neutrinos affected by in-vacuo dispersion. This is due to the size of the time window that one must adopt in order to investigate the hypothesis of in-vacuo dispersion. Standard searches of GRB neutrinos (assuming $\eta=0$, no in-vacuo dispersion) can assume that the neutrino and the electromagnetic signal from the GRB travel essentially at the same speed, and therefore their GRB-neutrino candidates are looked for within a tiny time window, rendering background issues insignificant. For $\eta\neq 0$, also taking into account that the neutrino energy will have some uncertainty (and however one will inevitably always end up testing an hypothesis of $\eta$ taking values in a certain range), the difference in observation time between the neutrino and the electromagnetic signal would be sizably uncertain, and the time window for searches of GRB-neutrino candidates would have to be correspondingly large, resulting in background issues such that one could not be sure that a specific GRB-neutrino association is correct.
Evidently the effectiveness of such a statistical approach depends rather crucially on the criteria used for the selection of candidate GRB neutrinos. And our main objective here is to test a novel proposal for these criteria. Testing alternative strategies of analysis on presently-available IceCube data is also important because it can set the stage for a later, more mature, phase of the research program, eventually using also data from KM3Net [35] and IceCube-Gen2 [36].
Previous searches [21, 22, 23, 24, 20, 33, 34] of GRB neutrinos affected by in-vacuo dispersion looked for neutrinos within a time window of fixed size (neutrino-energy-independent) after the GRB onset, and put on the same footing both GRBs whose redshift has not been measured (then having to estimate that redshift crudely, since the conjectured effect is redshift dependent) and GRBs of known redshift. Including GRBs whose redshift has not been measured has the advantage of a larger number of candidate GRB neutrinos, but of course renders the analysis vulnerable to the assumptions made to roughly estimate the redshifts. Moreover, the fixed-size time window, while easily handled computationally, imposes a restriction of the analysis to a corresponding limited range of neutrino energies: since the sought effect grows linearly with energy, a time window adapted to energies of, say, 100 TeV will inevitably be too small for neutrinos with energy much greater than 100 TeV, and in an appropriate sense it would also be too large a time window for neutrinos of energy smaller than 100 TeV (when the time window is much wider than the one really needed by the sought effect, the analysis ends up being dominated by the background). In spite of these limitations, Ref. [34], the latest such study (of which some of us were authors), provided the estimate $\eta=21.7\pm 4.5$ with a $p$-value [^2] As will be clearer for our readers when, later in this manuscript, we derive an analogous $p$-value for our novel selection criteria, this $p$-value, which in Ref. [34] was called false alarm probability, is the probability to find at least as many (and as good) GRB-neutrino candidates as were actually found in the data under the null hypothesis of no in-vacuo dispersion ($\eta=0$), whereas $\eta=21.7\pm 4.5$ is an estimate of $\eta$ obtained from the actually found GRB-neutrino candidates. $P_{[34]}=0.007$.
We here take the results of Ref. [34] as a starting point for testing a novel strategy of analysis. Rather than adopting a fixed-size time window and correspondingly restricting the neutrino-energy range, we consider all the neutrinos in the sample, regardless of the energy, but restrict the GRB catalogue to GRBs of known redshift. Any GRB and neutrino times of arrival are then regarded as compatible if they satisfy Eq. (1) for some $\eta\in\left[12.7,30.7\right]$ (the two-standard-deviation interval obtained from the estimate $\eta=21.7\pm 4.5$ found in Ref. [34]) within two standard uncertainties in the neutrino energy. As here shown in Sec. IV, the fact that, by relying exclusively on GRBs of known redshift, we find fewer GRB-neutrino candidates is more than compensated, for what concerns the overall statistical significance, by the sharper setup of the analysis which is available when the redshift of the GRBs is known.
[…]
II IceCube neutrinos and GRBs of known redshift
Our analysis relies on the latest HESE data release by IceCube (7 September 2023), available at [39]. We chose to work with the HESE sample because it is the IceCube sample least contaminated by atmospheric neutrinos, especially above 50-60 TeV [40]. As done in previous analogous studies [21, 22, 23, 24, 20, 34], we restrict to shower (or cascade) events, since track events have very poor energy estimates (and of course the quality of the energy information is crucial for in-vacuo-dispersion studies) [41, 42]. An additional advantage of working with shower events is that they are even more robust against atmospheric contamination (for the shower morphology, the atmospheric neutrino flux above 1 TeV is suppressed by about an order of magnitude with respect to the track morphology [43]).
A sizable part of the effort we devoted to this study was aimed at compiling a reliable catalogue of GRBs with actually measured (and not just estimated or loosely constrained) redshifts. In fact, we found that all already available GRB catalogues either imposed further constraints on the GRB properties (narrower observation window, restriction to specific GRB observatories, GRB type, etc.) or were not sufficiently accurate (misreporting or omitting direction or redshift data, not taking into account the latest relevant GCNs, etc.). The resulting list is reported here in Appendix A. In about 90% of cases it agrees with an analogous general-purpose catalogue, not restricted to known-redshift GRBs, maintained by the IceCube collaboration [44]. The previous study [34] relied on this resource, but, after checking one-by-one each of its known-redshift entries, we noticed that some of the information relevant for our analysis was inaccurately reported, probably due to the automated population of the database. In most cases this was quickly corrected according to the web catalogue [45] maintained by Jochen Greiner at the MPE, which we found to be the most accurate and comprehensive among publicly available lists. Checking by hand all the relevant GCNs and consulting additional catalogues of known-redshift GRBs, we were then able to correct the few errors left and also identify a few known-redshift GRBs whose redshift is not reported in Greiner’s table (see Appendix A for details).
In the end, all these efforts did not actually pay off in terms of a significant improvement of our analysis: as regards the four GRBs of known redshift playing a crucial role in our main analysis (see Sec. IV), our corrected and augmented catalogue agrees with both [44] and [45]. Still, we hope that the effort we put in preparing Appendix A will be of service to the community. In particular, our Appendix A should represent a tangible improvement over both [44] and [45] for researchers exploring alternative search strategies for GRB neutrinos affected by in-vacuo dispersion.
III Simulated data for the statistical-significance analysis and background estimate
As stressed above, the in-vacuo-dispersion significance of our findings can only be investigated using a statistical approach based on the $p$-value, i.e. estimating numerically the probability that one could find GRB-neutrino candidates in such good agreement with the in-vacuo-dispersion hypothesis just by accident, as an outlier of the type of findings one would expect within the null hypothesis $\eta=0$ of no in-vacuo dispersion.
As done in the previous neutrino studies [^3] The interested reader can find applications of this method of statistical analysis to studies not involving neutrinos, e.g., in Refs. [38, 46, 47, 48, 49, 50]. of Refs. [21, 22, 23], we do this by producing simulated data through transformations of the real data which reliably preserve their morphology while washing away any possible correlations of the type predicted by Eq. (1) between the time of arrival, the neutrino energy, and $D(z)$.
Our simulated data are obtained from the real data performing the following independent manipulations:
$\bullet$ we act on the neutrino observation times with a random permutation and a random periodic time translation (the periodicity makes it sure that they stay within the actual IceCube observation window);
$\bullet$ we act on the neutrino directions with a random rotation around the Earth’s axis;
$\bullet$ we act on the GRB directions with a random permutation and a random rotation around the Galactic axis.
A combination of these manipulations is arguably the most general transformation of the true data not affecting their morphology in any relevant respect. In fact:
$\bullet$ the neutrino observation times are statistically compatible with their being uniformly distributed within the IceCube observation window, and there are no a-priori reason for expecting a change in the HESE neutrino rate in this period;
$\bullet$ the efficiency of the IceCube detector, whose axis is closely aligned to the Earth’s axis, is virtually independent of the right ascension, but not of the declination, of incoming neutrinos, as clearly evidenced in the neutrino angular data; we choose not to reshuffle the neutrino directions so as not to spoil their evident correlation with the neutrino energies (the higher the neutrino energy, the more accurate its reconstructed direction);
$\bullet$ the directions of known-redshift GRBs are independent of galactic longitude, but are correlated to some extent with galactic latitude, as the dust lying around the galactic plane makes it much more difficult to determine the redshift of a GRB in that region; also, we do not act on the GRB observation times because we find that the detection rate of known-redshift GRBs is declining with time, probably another selection effect due to older events having been more likely followed by the long-term, host-galaxy observations which are often needed to determine the redshift of a GRB.
As a first application of our simulated data we can estimate the overall background that is expected with the criteria adopted by our strategy of analysis. For this purpose, we generated $10^{5}$ instances of simulated data [^4] In the following, whenever we refer to simulated data, it is understood that we did all computations based on $10^{5}$ instances. and we used them to estimate the expected number of HESE shower neutrinos that in the absence of in-vacuo dispersion could accidentally be associated with a GRB of known redshift according to the time-energy and direction compatibility conditions specified above (with $\eta\in[12.7,30.7]$ and $\mathcal{S}_{cut}=1/4\pi$). We find that, on average, our criteria should pick up 1.02 such accidental GRB-neutrino associations.
IV Main analysis
Equipped with the preparatory work reported in the previous sections, we are now ready to discuss our main analysis. We begin by looking for GRB-neutrino candidates, using the data described in Sec. II. For each GRB-neutrino pair selected by our directional criteria ($\mathcal{S}\geq 1/4\pi$) we describe the $\Delta t$ of Eq. (1) as the difference between the neutrino observation time [^5] If there was no in-vacuo dispersion, a GRB neutrino would be observed (nearly-)simultaneously with the GRB photons. We attribute the whole of the time-of-arrival difference to the $\Delta t$ of the neutrino, since photons observed from GRBs are of much lower energies than our neutrinos and the effect we are studying depends linearly on energy. Some quantum-spacetime scenarios predict in-vacuo dispersion with different magnitude for photons and neutrinos (and if the magnitude was much stronger for photons this assumption of our analysis would not be satisfied); it is however noteworthy that effective-field-theory approaches to quantum properties of spacetime predict that the type of pure in-vacuo dispersion here studied is actually forbidden for photons whereas it is allowed for neutrinos [17, 51]. and the observation time of the GRB, which we label as $\Delta t_{candidate}$, and then we use Eq. (1) and the known redshift of the GRB, which we label as $z_{GRB}$, to convert the requirement that $\eta$ should be within the interval $[12.7,30.7]$ into a corresponding range of allowed values for the neutrino energy. We then consider the GRB-neutrino pair a GRB-neutrino candidate if the energy of the neutrino is compatible with that energy range within two standard deviations (assuming 10% uncertainty in the energy of the neutrino [41]). We find that there are four such GRB-neutrino candidates. Fig. 1 provides a visual characterization of the fact that the properties of these four GRB-neutrino candidates match rather well the expectations of the in-vacuo-dispersion scenario of Eq. (1).
Figure 1: The values of $E$ and $\Delta t^{*}$ of our four GRB-neutrino candidates line up rather nicely according to the expectations of in-vacuo dispersion. Only one of them, here shown as a violet square, was taken into account also in the analysis of Ref. [34]. The content of this figure should be assessed also keeping in mind the results here reported in Sec. III, suggesting that it is likely that one of our four GRB-neutrino candidates is background. Based on the estimate $\eta=21.7\pm 4.5$ reported in Ref. [34], the dark-gray line corresponds to $\eta=21.7$, while the light-gray band highlights our region of interest in the $E$-$\Delta t^{*}$ plane, corresponding to $\eta\in[12.7,30.7]$ via Eq. (1) (see Sec. I for comments on this interval).
In Fig. 1, $\Delta t^{*}$ is defined as
$\Delta t^{*}\equiv{D(1)\over D(z_{GRB})}\Delta t_{candidate}$
(the numerical factor $D(1)$ is introduced only for the convenience of having a $\Delta t^{*}$ with dimensions of time), and according to Eq. (1) one should find a linear relationship between $E$ and $\Delta t^{*}$,
$\Delta t^{*}=\eta D(1)\frac{E}{M_{P}}\,.$
It is clearly encouraging for the hypothesis of in-vacuo dispersion that our criteria find four GRB-neutrino candidates (some details on these four GRB-neutrino candidates are provided in Appendix B), whereas according to the pure-background null hypothesis we expected to typically find only one (see Sec. III). Further encouragement comes from the fact that in Fig. 1 our four GRB-neutrino candidates line up rather nicely, as measured by their correlation coefficient $r(\Delta t^{*},E)=0.9985$. In particular, three of the four GRB-neutrino candidates are compatible with a remarkably narrow range of values of $\eta$, an observation which strengthens the appeal of the conjecture that three of our four GRB-neutrino candidates actually might be GRB neutrinos affected by in-vacuo dispersion.
The fact that our findings reported in Fig. 1 fit very well with the estimate $\eta=21.7\pm 4.5$, which resulted from the analysis of Ref. [34], is noteworthy since, because of the different criteria adopted for the selection of GRB-neutrino candidates, the two analyses only share one GRB-neutrino candidate, the one shown as a violet square in Fig. 1. The analysis leading to the estimate $\eta=21.7\pm 4.5$, in Ref. [34] ended up focusing on seven GRB-neutrino candidates, but only one of them involved a GRB with measured redshift.
It is of course necessary to quantify statistically this observation that the data we are analyzing match well the expectations of in-vacuo dispersion. For this purpose we rely on the same strategy of characterization of statistical significance adopted in Ref. [34]. In parts of the next Section we shall contemplate alternative strategies of characterization of statistical significance, but the main objective of the study we are here reporting is to compare the GRB-neutrino-candidate selection criteria here adopted with the selection criteria adopted in Ref. [34], and this comparison is of course more transparent if the two selection strategies are assessed using the same characterization of statistical significance. The main analysis of Ref. [34] focused on seven GRB-neutrino candidates and established how frequently simulated data produced at least seven GRB-neutrino candidates with correlation at least as high as the correlation of the GRB-neutrino candidates found in the true data. This happened in about 0.7% of the trials, yielding the above-mentioned $p$-value $P_{[34]}=0.007$. Proceeding in an analogous way with the selection criteria here adopted, we start by using our simulated data (Sec. III) to estimate how frequently the null hypothesis of no in-vacuo dispersion ($\eta=0$) would produce at least four GRB-neutrino candidates according to our new selection criteria, finding a $p$-value $P_{n}=0.019$. We then proceed to estimate how frequently the null hypothesis would accidentally exhibit at least four GRB-neutrino candidates with correlation $r(\Delta t^{*},E)\geq 0.9985$, finding a $p$-value $P_{r}=0.006$ (which corresponds to a $2.8\sigma$ significance in Gaussian statistics).
V Avenues for refining the strategy of analysis
We consider as our main result the $p$-value $P_{r}=0.006$ derived in the previous Section, which, while being still far from conclusive, testifies to the discovery potential of the novel GRB-neutrino-candidate selection criteria here advocated, which was our main objective.
In this Section we discuss and investigate some observations which could be used to refine the strategy of analysis. We believe that some of the observations reported in this Section deserve being considered in setting up in-vacuo-dispersion analyses of future neutrino data, even though some of the quantitative aspects of this Section should be assessed with caution since they were produced in a second phase of our investigations after seeing the data, and might therefore be affected by the sources of bias that are well known for unblind analyses. In particular, some of the observations reported in this Section are at least in part inspired by the fact that we saw that among our four GRB-neutrino candidates there are three that are described particularly well according to Eq. (1), and by our perception that the procedure used in Sec. IV to estimate the significance of our findings did not “benefit” sufficiently from the noteworthy properties of those three best GRB-neutrino candidates.
V.1 Alternative choices of $S_{cut}$
In planning our study we ended up adopting the rather prudent criterion of directional consistency characterized by the intuitively natural choice $\mathcal{S}_{cut}=1/4\pi$ (see Sec. I). After completing our main analysis we perceived the need to quantify how frequently this criterion was picking up accidental directional associations between a neutrino and a GRB. Using our simulated data, we checked that the choice $\mathcal{S}_{cut}=1/4\pi$ entails a 7.8% probability of accidental directional association between a neutrino and a GRB, which is of course partly responsible for the (large but) tolerable expectation that about one background GRB-neutrino candidate should be accidentally picked up by our overall selection criteria (see Sec. III).
We then checked how strongly the $p$-value $P_{r}$, which we regard as our main result, depends on the choice of $\mathcal{S}_{cut}$. Fig. 2 shows the dependence of $P_{r}$ on $\mathcal{S}_{cut}$ for $\mathcal{S}_{cut}\in[0.032,0.241]$. The rationale for exploring this particular range of values is that, as checked using our simulated data, $\mathcal{S}_{cut}=0.241$ entails a 4.5% probability of accidental directional association, the nominal “two standard deviation” reference, whereas $\mathcal{S}_{cut}=0.032$ corresponds to a 10% probability of spurious association. As clearly evidenced by Fig. 2, our main result $P_{r}$ does not depend strongly on the choice of $\mathcal{S}_{cut}$. It is potentially intriguing that for values of $\mathcal{S}_{cut}$ just below $1/4\pi$ we would have found an even smaller $P_{r}$, but, in keeping with the considerations of Sec. I, we currently regard this circumstance as a mere numerical accident occurring with the presently-available data (rather than a hint for optimizing the analysis of future data).
Figure 2: The $p$-value $P_{r}$ (see Sec. IV for the definition) as a function of $\mathcal{S}_{cut}$. The blue square corresponds to the choice $\mathcal{S}_{cut}=1/4\pi$ of Sec. IV.
V.2 Improved correlation
One challenge whose handling might require further methodological refinements in future studies concerns the choice of an appropriate statistic quantifying the agreement between the properties of the found GRB-neutrino candidates and Eq. (1). Investigations of this aspect should take into account the fact that the time window for in-vacuo dispersion searches of GRB neutrinos will always be rather large, even if, as one might hope, the $\eta$ range on which one would focus could become narrower with the accrual of more data. A relatively small uncertainty on $\eta$ still translates (especially at high neutrino energies and high redshift) into a rather large time window, large enough for a tangible probability of background contamination of the GRB-neutrino candidates. These concerns are even more severe if one allows for non-systematic, fuzzy in-vacuo dispersion effects of the type mentioned here briefly in Sec. I (and discussed in more detail in, e.g., Refs. [18, 19, 26]), since in that case searches of GRB-neutrino candidates would always have to explore a range of $\eta$ with width at least $\delta\eta$.
Awareness of these issues should encourage investigations aimed at quantifying the agreement between the found GRB-neutrino candidates and the relevant in-vacuo-dispersion models in ways that are both robust against contamination by the background and capable of profiting from all the implications of in-vacuo-dispersion.
In our main analysis, reported in Sec. IV, we followed Ref. [34] and characterized the agreement between our findings and Eq. (1) in terms of the standard (Pearson) correlation coefficient $r(\Delta t^{*},E)$. However, $r$, while giving an appropriate weight to the linear relationship between $E$ and $\Delta t^{*}$ predicted by Eq. (1), is completely insensitive to the fact that, according to Eq. (1), one should have $\Delta t^{*}\rightarrow 0$ for $E\rightarrow 0$. We observe that this shortcoming can be remedied by replacing the sample covariance and the sample variances appearing in the definition of $r$ with the corresponding raw (non-central) sample moments, i.e., by replacing $r(\Delta t^{*},E)$ with the corresponding non-central correlation coefficient
$r_{0}(\Delta t^{*},E)=\frac{\sum_{i}\Delta t^{*}_{i}E_{i}}{\sqrt{\sum_{j}(\Delta t^{*}_{j})^{2}\sum_{k}E_{k}^{2}}}.$
It is clear that $r_{0}$ indeed rewards GRB-neutrino candidates fitting the expectation that $E$ depends linearly on $\Delta t^{*}$ with negligible intercept.
We can test the effectiveness of this improved statistic by replacing $r$ with $r_{0}$ in our main analysis (Sec. IV). Our four GRB-neutrino candidates have $r_{0}(\Delta t^{*},E)=0.9989$, and requiring that our simulated data (Sec. III) accidentally produce at least four GRB-neutrino candidates with $r_{0}(\Delta t^{*},E)\geq 0.9989$, we find a $p$-value $P_{r_{0}}=0.004$.
The fact that $P_{r_{0}}<P_{r}$ suggests that our GRB-neutrino candidates are indeed consistent with the expectation that $\Delta t^{*}\rightarrow 0$ for $E\rightarrow 0$.
V.3 Taking into account the background estimate
Regardless of the statistic used to quantify the agreement of the found GRB-neutrino candidates with Eq. (1), one could also try to tame background contamination by taking explicitly into account its contribution before computing the statistic.
Looking at the GRB-neutrino candidate with $E=302$ TeV in Fig. 1, one can easily understand these concerns: even though we expect that one of our four GRB-neutrino candidates should be background, in our main analysis (following Ref. [34]) we computed the correlation coefficient of all of them. Future studies might thus contemplate the possibility of taking into account the background estimate. More explicitly, if one finds $N$ GRB-neutrino candidates and estimates that there is a large probability, e.g., 90%, that at least $M$ should be background, it may be appropriate to assess the agreement between the found GRB-neutrino candidates and Eq. (1) using only the best $N-M$ candidates.
In the case of the study here reported, the probability of finding at least one background candidate is still too low (about 60%) to justify disregarding our worst candidate. Nevertheless, for mere illustrative purposes, we briefly discuss how to amend our main analysis if one of our four candidates could be actually ignored. First, we compute the correlation coefficients of all possible choices of three of our four GRB-neutrino candidates and find that the highest correlation is $r(\Delta t^{*},E)=0.999997$ (found indeed excluding the candidate with $E=302$ TeV). We then use our simulated data to estimate how frequently the null hypothesis would accidentally produce at least four GRB-neutrino candidates such that three of them have correlation $r(\Delta t^{*},E)\geq 0.999997$, finding a $p$-value $P_{rb}=0.002$.
V.4 Scanning a range of values of $\eta$
The fact that we are restricting our analysis to GRBs with measured redshift allows us to explore another statistic, which we expect to be less vulnerable than correlation to background contamination.
For simplicity we illustrate and discuss the strategy of analysis based on this alternative statistic applying it directly to the data already analyzed in this paper. For every given value of $\eta$ in our interval of interest $[12.7,30.7]$, we find the number $N(\eta)$ of GRB-neutrino candidates which are directionally compatible in the usual sense ($\mathcal{S}\geq 1/4\pi$) and satisfy Eq. (1) for that specific value of $\eta$ within two standard uncertainties in the neutrino energy. We then use simulated data to estimate the probability that the null hypothesis would accidentally produce at least $N(\eta)$ such GRB-neutrino candidates, finding a corresponding (local) $p$-value $p(\eta)$. The resulting “$p$-curve” is reproduced here as Fig. 3.
Figure 3: The $p$-curve for $\eta\in[12.7,30.7]$. It is obtained estimating how frequently simulated data produce at least the same number of GRB-neutrino candidates as the real data for each value of $\eta$ (see text for details). The horizontal dashed line is drawn at $p=0.05$.
Consistently with the fact, already visible in Fig. 1, that our four GRB-neutrino candidates are remarkably compatible with a narrow range of values of $\eta$, we see that the $p$-curve of Fig. 3 shows a deep valley surrounded by considerably higher values of $p$. In particular, $p(\eta)$ is less than 0.05 (marked as a dashed horizontal line in Fig. 3) for $\eta\in[18.3,25.1]$. The minimum of the $p$-curve is found at $\eta={20.8}$, fully consistent with the estimate $21.7\pm 4.5$ reported in Ref. [34].
The statistical significance of our $p$-curve can be characterized estimating how frequently the minimum of the analogous $p$-curves built out of simulated data happens to be less than or equal to the minimum value $p(20.8)=0.004$ attained in Fig. 3. The resulting (global) $p$-value is $P_{ns}=0.012$.
The fact that $P_{ns}$ and $P_{r}$, resulting from considerably different strategies of analysis, are of the same order of magnitude suggests that both values are in fact genuine characterizations of the statistical significance of our results, i.e., that they are not too dependent on the specific properties of the corresponding statistics.
V.5 Exploring $\eta$ outside the range $[12.7,30.7]$
In our main analysis we let $\eta$ vary in the range $[12.7,30.7]$ favored by the previous study [34], which adopted different selection criteria, obtaining results in remarkable agreement with those of Ref. [34]. Still, it is interesting to check whether our new GRB-neutrino-candidate selection criteria would have yielded even stronger results for values of $\eta$ outside the range $[12.7,30.7]$. A natural way of doing this is to extend the domain of the $p$-curve introduced in the previous Subsection. In Fig. 4 we show the $p$-curve for $\eta\in[-50,50]$.
Figure 4: The same $p$-curve of Fig. 3 for $\eta\in[-50,50]$.
Looking at Fig. 4, it is clear that also our new GRB-neutrino-candidate selection criteria favor the range $[12.7,30.7]$ for $\eta$. The fact that we find no other values of $\eta$ with comparable $p(\eta)$ is also potentially significant for a much-studied effective-field-theory description of quantum-spacetime properties [52, 51], which in principle allows independent in-vacuo-dispersion parameters for the two neutrino helicities. The prevalent theoretical prejudice is that the two helicities should have the same in-vacuo-dispersion parameter, but it has also been argued that the two helicities might have $\eta$ with the same modulus but opposite sign [21, 24]. The $p$-curve reproduced in Fig. 4 clearly does not provide any encouragement for this helicity-dependence scenario.
VI Outlook
The essential equivalence of the $p$-values $P_{[34]}$, characterizing the statistical significance of the previous analysis [34], and $P_{r}$, found here in Sec. IV and characterizing in the same way the significance of our current analysis, shows that the new, redshift-based selection criteria for GRB-neutrino candidates proposed in Sec. I are competitive with the ones adopted in Ref. [34]. This means that the improvement in data quality provided by the knowledge of the GRB redshifts is more than sufficient to compensate for the loss of most of the GRB dataset (we measure the redshift of about 1/10 of observed GRBs). We feel that this is our main result, as it effectively opens up a new avenue, complementary to the approach pursued in Ref. [34], to investigate in-vacuo dispersion for neutrinos.
Considering the very significant differences in their selection criteria (as stressed above, only one GRB-neutrino candidate figures in both analyses), it is noteworthy that the analysis here reported and the one reported in Ref. [34] both favor the same range of values of $\eta$. This might suggest that, if one were somehow able to combine the two analyses, then the overall evidence in favor of in-vacuo dispersion provided by presently-available IceCube data might be stronger than indicated by the individual statistical significance of each analysis. However, we do not yet see how such a combination of selection criteria could be achieved in a logically consistent manner, mostly because sharp knowledge of the GRB redshifts is a structural feature of the strategy of analysis advocated here.
We expect that the redshift-based strategy articulated in the previous sections will prove to be the best starting point for any future refinements, but we do not exclude that a suitable way may still be found for including GRBs whose redshift has not been measured (of course, from the perspective of our approach, somehow estimating the redshift of those GRBs with a suitably large uncertainty). Such a refinement of our approach could prove even more valuable in a few years, as more IceCube and KM3NeT data is accrued.
At least equally important for the analysis of future IceCube and KM3NeT data would be to remove the other key limitation of our selection criteria, the restriction to shower-event neutrinos. This could in principle be achieved by incorporating into the analysis the known [42] large and non-Gaussian probability distribution (whose peak lies sizably above the visible energy) that relates the primary neutrino energy to the visible energy in track events.
Acknowledgements.
We thank Anastasia Tsvetkova for guidance on the available GRB catalogues and resources, which was very valuable in the early stages of our work on Appendix A. G.A.-C. and G.G. are grateful for financial support by the Programme STAR Plus, funded by Federico II University and Compagnia di San Paolo, and by the MIUR, PRIN 2017 grant 20179ZF5KS. G.R.’s work on this project was supported by the National Science Centre grant 2019/33/B/ST2/00050. G.D.’s work on this project was supported by the Research Council of Norway, project number 301718, and by the Beatriu de Pinós programme (BP 2023). G.F.’s work on this project was supported by “Fondazione Angelo Della Riccia” and “The Foundation Blanceflor”. This work also falls within the scopes of the COST Action CA23130 “Bridging high and low energies in search of quantum gravity” and the COST Action CA18108 ”Quantum Gravity phenomenology in the Multi-Messenger era”.
Appendix A Known-redshift GRB data
The following table contains the GRB data used in our analysis. We include all GRBs observed up to December 2022 for which a definite redshift determination is available in the literature. We did not include GRBs with only upper and/or lower bounds on redshift. For the selection of the sample and the redshift values $z$ we relied mostly on Greiner’s catalogue [45], referred to in our list as MP. Searching the literature for other GRBs whose redshift was measured, we found 45 which are not acknowledged as such in the summary table of [45]. For these GRBs we provide references to the relevant works and GCNs in the column “$z$ ref.” of our table.
The angular data and the trigger times (TT) of the GRBs in our sample were downloaded from the web catalogue maintained by the IceCube collaboration [44], and were then manually checked against the primary GCN sources referred to in Greiner’s list [45]. If the right ascension (RA) or the declination (decl) values reported by [44] were no more than 0.001° off the best values we found in [45], we did not correct them, as our analysis is completely insensitive to errors of that order of magnitude. For the same reason, we do not provide the uncertainties affecting the angular positions, since we checked that they are all less than 0.001°. Accordingly, we report our angular values with three digital figures, with the understanding that the last one is affected by an uncertainty of order 0.001°.
Table 1: List of GRBs with measured redshift
     Name RA (°) decl (°) TT (MJD) $z$ $z$ ref.
GRB221226B 22.909 -41.527 59939.945 2.694 MP
GRB221110A 29.100 -27.294 59893.103 4.06 MP
GRB221009A 288.264 19.773 59861.553 0.151 MP
GRB221006A 337.246 15.714 59858.040 0.731 [53, 54]
GRB220813A 81.532 -33.016 59804.808 0.82 [55]
GRB220627A 201.369 -32.426 59757.890 3.084 MP
GRB220611A 66.515 -37.260 59741.751 2.3608 MP
GRB220527A 323.528 -14.972 59726.387 0.857 MP
GRB220521A 275.230 10.372 59720.972 5.6 MP
GRB220219B 240.914 31.234 59629.394 0.293 MP
GRB220117A 91.572 -28.437 59596.680 4.961 MP
GRB220107A 169.807 34.171 59586.615 1.246 MP
GRB220101A 1.353 31.769 59580.215 4.618 MP
GRB211227A 132.149 -2.735 59575.981 0.228 [56, 57]
GRB211211A 212.292 27.889 59559.549 0.073 [58, 59]
GRB211207A 149.624 -24.359 59555.870 2.272 MP
GRB211024B 154.713 24.568 59511.931 1.1137 MP
GRB211023B 170.310 39.136 59510.879 0.862 MP
GRB211023A 73.154 85.324 59510.546 0.39 MP
GRB210919A 80.254 1.312 59476.020 0.2411 MP
GRB210905A 309.048 -44.440 59462.009 6.318 MP
GRB210822A 304.438 5.283 59448.388 1.736 MP
GRB210731A 300.305 -28.061 59426.931 1.2525 MP
GRB210722A 27.030 -6.347 59417.871 1.145 MP
GRB210702A 168.578 -36.747 59397.797 1.16 MP
GRB210619B 319.718 33.850 59385.000 1.937 MP
GRB210610B 243.918 14.399 59375.827 1.13 MP
GRB210610A 204.282 14.465 59375.628 3.54 MP
GRB210517A 358.224 -39.102 59351.228 2.486 MP
GRB210504A 222.392 -30.534 59338.580 2.077 MP
GRB210420B 254.325 42.570 59324.774 1.4 MP
GRB210411C 296.612 -39.398 59315.629 2.826 MP
GRB210323A 317.947 25.369 59296.918 0.733 MP
GRB210321A 87.895 70.130 59294.301 1.487 MP
GRB210312B 155.814 76.869 59285.870 1.069 MP
GRB210222B 154.606 -14.932 59267.943 2.198 MP
GRB210210A 262.771 14.663 59255.084 0.715 MP
GRB210204A 117.081 11.410 59249.270 0.876 MP
GRB210116A 123.814 -5.867 59230.246 2.514 MP
GRB210112A 219.006 33.054 59226.067 2 [60, 61]
GRB210104A 103.772 64.676 59218.477 0.46 MP
GRB201221D 171.059 42.144 59204.963 1.045 MP
GRB201221A 214.480 -45.416 59204.298 5.7 MP
GRB201216C 16.370 16.516 59199.963 1.1 MP
GRB201104B 5.215 7.842 59157.732 1.954 MP
GRB201103B 42.185 12.137 59156.755 1.105 MP
GRB201024A 125.952 3.354 59146.117 0.999 MP
GRB201021C 12.529 -55.866 59143.852 1.07 MP
GRB201020B 75.470 77.068 59142.732 0.804 MP
GRB201020A 261.228 31.428 59142.241 2.903 MP
GRB201015A 354.319 53.416 59137.952 0.426 MP
GRB201014A 20.796 27.660 59136.950 4.56 MP
GRB200829A 251.205 72.329 59090.582 1.25 MP
GRB200826A 6.786 34.027 59087.187 0.7481 MP
GRB200613A 153.042 45.754 59013.229 1.227 MP
GRB200524A 213.043 60.905 58993.211 1.256 MP
GRB200522A 5.682 -0.283 58991.487 0.554 MP
GRB200411A 47.664 -52.318 58950.187 0.7 MP
GRB200219A 342.638 -59.120 58898.317 0.48 MP
GRB200205B 107.788 -56.488 58884.807 1.465 MP
GRB191221B 154.830 -38.158 58838.861 1.148 MP
GRB191031D 283.289 47.644 58787.891 0.5 MP
GRB191019A 340.025 -17.328 58775.634 0.248 MP
GRB191011A 44.728 -27.845 58767.192 1.722 MP
GRB191004B 49.205 -39.634 58760.898 3.503 MP
GRB190919B 311.877 -44.695 58745.991 3.225 MP
GRB190829A $4$8.958 58724.830 0.0785 MP
GRB190719C 240.207 13.000 58683.624 2.469 MP
GRB190627A 244.828 -5.289 58661.471 1.942 MP
GRB190613A 182.529 67.235 58647.172 2.78 MP
GRB190324A 49.616 -47.215 58566.947 1.1715 MP
GRB190114C 54.505 -26.946 58497.873 0.425 MP
GRB190114A 65.544 2.192 58497.133 3.3765 MP
GRB190106A 29.880 23.846 58489.566 1.859 MP
GRB181201A 319.297 -12.631 58453.110 0.45 [62, 63]
GRB181123B 184.367 14.598 58445.231 1.754 MP
GRB181110A 302.318 -36.897 58432.364 1.505 MP
GRB181020A 13.982 -47.381 58411.792 2.938 MP
GRB181010A 52.570 -23.038 58401.247 1.39 MP
GRB180914B 332.356 25.062 58375.766 1.096 MP
GRB180805B 25.782 -17.494 58335.543 0.661 MP
GRB180728A 253.565 -54.044 58327.728 0.117 MP
GRB180727A 346.666 -63.052 58326.594 2 MP
GRB180720B 0.529 -2.919 58319.598 0.654 MP
GRB180703A 6.468 -67.179 58302.876 0.6678 MP
GRB180624A 318.098 -2.338 58293.576 2.855 MP
GRB180620B 357.521 -57.962 58289.660 1.1175 MP
GRB180618A 169.941 73.837 58287.030 0.544 MP
GRB180510B 77.969 -62.324 58248.844 1.305 MP
GRB180418A 170.122 24.933 58226.281 1.55 MP
GRB180404A 83.548 -37.168 58212.032 1 MP
GRB180329B 82.904 -23.690 58206.589 1.998 MP
GRB180325A 157.428 24.464 58202.078 2.248 MP
GRB180314A 99.265 -24.496 58191.030 1.445 MP
GRB180205A 126.820 11.542 58154.184 1.409 MP
GRB180115A 12.039 -15.630 58133.178 2.487 MP
GRB171222A 148.278 35.627 58109.684 2.409 MP
GRB171205A 167.415 -12.588 58092.306 0.0368 MP
GRB171020A 39.248 15.204 58046.963 1.87 MP
GRB171010A 66.581 -10.463 58036.792 0.3293 MP
GRB170903A 254.526 34.979 57999.534 0.886 MP
GRB170817A 197.450 -23.381 57982.529 0.0098 MP
GRB170728B 237.981 70.122 57962.961 1.27 MP
GRB170728A 58.888 12.182 57962.287 1.49 MP
GRB170714A 34.350 1.991 57948.518 0.793 MP
GRB170705A 191.704 18.307 57939.115 2.01 MP
GRB170607A 7.366 9.243 57911.971 0.557 MP
GRB170604A 342.656 -15.412 57908.798 1.329 MP
GRB170531B 286.884 -16.418 57904.918 2.366 MP
GRB170519A 163.427 25.374 57892.215 0.818 MP
GRB170428A 330.078 26.916 57871.384 0.454 MP
GRB170405A 219.828 -25.243 57848.777 3.51 MP
GRB170214A 256.341 -1.888 57798.649 2.53 MP
GRB170202A 152.514 5.012 57786.769 3.645 MP
GRB170127B 19.977 -30.358 57780.634 2.2 MP
GRB170113A 61.733 -71.943 57766.420 1.968 MP
GRB161219B 91.714 -26.792 57741.784 0.1475 MP
GRB161129A 316.228 32.135 57721.300 0.645 MP
GRB161117A 322.052 -29.614 57709.066 1.549 MP
GRB161108A 180.788 24.868 57700.148 1.159 MP
GRB161104A 77.894 -51.460 57696.404 0.79 MP
GRB161023A 311.022 -47.663 57684.944 2.708 MP
GRB161017A 142.769 43.127 57678.744 2.013 MP
GRB161014A 332.648 7.469 57675.522 2.823 MP
GRB161001A 71.920 -57.261 57662.045 0.67 MP
GRB160821B 279.977 62.392 57621.937 0.16 MP
GRB160804A 221.630 9.999 57604.064 0.736 MP
GRB160629A 4.863 76.967 57568.930 3.332 MP
GRB160625B 308.598 6.919 57564.945 1.406 MP
GRB160624A 330.193 29.644 57563.477 0.483 MP
GRB160623A 315.298 42.221 57562.208 0.367 MP
GRB160509A 311.754 76.108 57517.374 1.17 MP
GRB160425A 280.327 -54.360 57503.977 0.555 MP
GRB160410A 150.685 3.478 57488.215 1.717 MP
GRB160408A 122.625 71.128 57486.268 1.9 MP
GRB160327A 146.702 54.013 57474.386 5 MP
GRB160314A 112.790 17.000 57461.481 0.726 MP
GRB160303A 168.701 22.742 57450.455 1 MP
GRB160228A 107.316 26.932 57446.732 1.64 MP
GRB160227A 194.808 78.679 57445.814 2.38 MP
GRB160203A 161.951 -24.789 57421.092 3.52 MP
GRB160131A 78.168 -7.050 57418.348 0.971 MP
GRB160121A 109.088 -23.592 57408.577 1.96 MP
GRB160117B 132.195 -16.367 57404.583 0.87 MP
GRB151229A 329.370 -20.732 57385.285 1.4 MP
GRB151215A 93.584 35.516 57371.126 2.59 MP
GRB151112A 2.054 -61.663 57338.573 4.1 MP
GRB151111A 56.845 -44.161 57337.356 3.5 MP
GRB151031A 83.196 -39.122 57326.243 1.167 MP
GRB151029A 38.528 -35.386 57324.326 1.423 MP
GRB151027B 76.220 -6.450 57322.945 4.063 MP
GRB151027A 272.487 61.353 57322.166 0.81 MP
GRB151021A 337.644 -33.197 57316.062 2.33 MP
GRB150915A 319.658 -34.914 57280.888 1.968 MP
GRB150910A 5.667 33.473 57275.378 1.359 MP
GRB150831A 221.024 -25.635 57265.440 1.18 MP
GRB150821A 341.913 -57.894 57255.406 0.755 MP
GRB150818A 230.356 68.342 57252.484 0.282 MP
GRB150728A 292.229 33.916 57231.536 0.46 MP
GRB150727A 203.969 -18.325 57230.793 0.313 MP
GRB150616A 314.717 -53.394 57189.951 1.188 [64]
GRB150518A 234.201 16.330 57160.904 0.256 MP
GRB150514A 74.876 -60.968 57156.774 0.807 MP
GRB150424A 152.306 -26.631 57136.321 1 [65, 66]
GRB150423A 221.579 12.284 57135.269 1.394 MP
GRB150413A 190.425 71.841 57125.580 3.139 MP
GRB150403A 311.505 -62.711 57115.913 2.06 MP
GRB150323A 128.178 45.465 57104.118 0.593 MP
GRB150314A 126.670 63.834 57095.205 1.758 MP
GRB150301B 89.166 -57.970 57082.818 1.5169 MP
GRB150206A 10.074 -63.182 57059.604 2.087 MP
GRB150120B 39.291 8.078 57042.307 3.5 MP
GRB150120A 10.319 33.995 57042.123 0.46 MP
GRB150101B 188.020 -10.934 57023.641 0.134 MP
GRB141225A 138.779 33.792 57016.959 0.915 MP
GRB141221A 198.287 8.205 57012.338 1.452 MP
GRB141220A 195.066 32.146 57011.252 1.3195 MP
GRB141212A 39.125 18.147 57003.510 0.596 MP
GRB141121A 122.669 22.217 56982.150 1.47 MP
GRB141109A 144.531 -0.608 56970.243 2.993 MP
GRB141028A 322.602 -0.231 56958.455 2.33 MP
GRB141026A 44.084 26.928 56956.109 3.35 MP
GRB141004A 76.734 12.820 56934.973 0.573 MP
GRB140930B 6.348 24.295 56930.821 1.465 MP
GRB140907A 48.146 46.605 56907.672 1.21 MP
GRB140903A 238.014 27.603 56903.625 0.351 MP
GRB140808A 221.222 49.215 56877.037 3.29 MP
GRB140801A 44.069 30.938 56870.792 1.32 MP
GRB140713A 281.106 59.634 56851.780 0.935 MP
GRB140710A 41.068 35.499 56848.428 0.558 MP
GRB140703A 12.996 45.102 56841.026 3.14 MP
GRB140629A 248.977 41.877 56837.595 2.275 MP
GRB140623A 225.473 81.191 56831.223 1.92 MP
GRB140622A 317.173 -14.419 56830.400 0.959 MP
GRB140620A 281.871 49.731 56828.219 2.04 MP
GRB140614A 231.169 -79.129 56822.045 4.233 MP
GRB140606B 328.125 32.015 56814.133 0.384 MP
GRB140518A 227.252 42.418 56795.387 4.707 MP
GRB140515A 186.064 15.105 56792.384 6.32 MP
GRB140512A 289.370 -15.094 56789.814 0.725 MP
GRB140509A 46.594 -62.639 56786.099 2.4 MP
GRB140508A 255.466 46.780 56785.128 1.027 MP
GRB140506A 276.775 -55.636 56783.880 0.889 MP
GRB140430A 102.936 23.024 56777.857 1.6 MP
GRB140428A 194.368 28.385 56775.945 4.7 MP
GRB140423A 197.286 49.842 56770.355 3.26 MP
GRB140419A 126.990 46.240 56766.171 3.956 MP
GRB140331A 134.864 2.717 56747.243 1 [67]
GRB140318A 184.089 20.209 56734.006 1.02 MP
GRB140311A 209.305 0.642 56727.879 4.954 MP
GRB140304A 30.643 33.474 56720.557 5.283 MP
GRB140301A 69.558 -34.257 56717.642 1.416 MP
GRB140226A 221.492 14.993 56714.419 1.98 MP
GRB140213A 105.155 -73.137 56701.807 1.2076 MP
GRB140206A 145.334 66.761 56694.304 2.73 MP
GRB140129B 326.757 26.206 56686.536 0.43 MP
GRB140114A 188.522 27.951 56671.498 3 MP
GRB131231A 10.590 -1.653 56657.198 0.642 MP
GRB131229A 85.232 -4.396 56655.277 1 MP
GRB131227A 67.378 28.883 56653.198 5.3 MP
GRB131117A 332.331 -31.762 56613.024 4.042 MP
GRB131108A 156.502 9.662 56604.862 2.4 MP
GRB131105A 70.967 -62.995 56601.087 1.686 MP
GRB131103A 348.919 -44.640 56599.922 0.599 MP
GRB131030A 345.067 -5.368 56595.872 1.295 MP
GRB131011A 32.526 -4.411 56576.741 1.874 MP
GRB131004A 296.113 -2.958 56569.904 0.717 MP
GRB130925A 41.179 -26.153 56560.164 0.347 MP
GRB130907A 215.892 45.608 56542.902 1.238 MP
GRB130831A 358.624 29.430 56535.545 0.4791 MP
GRB130822A 27.922 -3.208 56526.663 0.154 MP
GRB130716A 179.574 63.053 56489.442 2.2 MP
GRB130702A 217.312 15.774 56475.004 0.145 MP
GRB130701A 357.229 36.100 56474.179 1.155 MP
GRB130615A 274.829 -68.161 56458.406 2.9 [68, 64]
GRB130612A 259.794 16.720 56455.141 2.006 MP
GRB130610A 224.420 28.207 56453.133 2.092 MP
GRB130606A 249.396 29.796 56449.878 5.913 MP
GRB130604A 250.187 68.226 56447.288 1.06 MP
GRB130603B 172.201 17.071 56446.659 0.356 MP
GRB130528A 139.505 87.301 56440.695 1.25 [69]
GRB130518A 355.668 47.465 56430.580 2.489 MP
GRB130515A 283.440 -54.279 56427.056 0.8 MP
GRB130514A 296.283 -7.976 56426.301 3.6 MP
GRB130511A 196.646 18.710 56423.480 1.3033 MP
GRB130505A 137.061 17.485 56417.349 2.27 MP
GRB130427B 314.898 -22.546 56409.556 2.78 MP
GRB130427A 173.137 27.699 56409.324 0.34 MP
GRB130420A 196.106 59.424 56402.311 1.297 MP
GRB130418A 149.037 13.667 56400.792 1.218 MP
GRB130408A 134.405 -32.361 56390.911 3.758 MP
GRB130215A 43.503 13.395 56338.063 0.597 MP
GRB130131B 173.956 15.038 56323.799 2.539 MP
GRB130131A 171.127 48.076 56323.581 1.55 MP
GRB121229A 190.101 -50.594 56290.209 2.707 MP
GRB121226A 168.642 -30.406 56287.798 1.37 MP
GRB121217A 153.710 -62.351 56278.303 3.1 [70]
GRB121211A 195.533 30.148 56272.574 1.023 MP
GRB121209A 326.787 -8.235 56270.916 2.1 MP
GRB121201A 13.467 -42.943 56262.518 3.385 MP
GRB121128A 300.600 54.300 56259.212 2.2 MP
GRB121123A 307.318 -11.860 56254.419 2.7 [71, 72]
GRB121027A 63.598 -58.830 56227.314 1.773 MP
GRB121024A 70.472 -12.291 56224.122 2.298 MP
GRB120923A 303.795 6.221 56193.220 7.8 MP
GRB120922A 234.748 -20.182 56192.938 3.1 MP
GRB120909A 275.736 -59.449 56179.070 3.93 MP
GRB120907A 74.750 -9.315 56177.017 0.97 MP
GRB120815A 273.958 -52.131 56154.093 2.358 MP
GRB120811C 199.683 62.301 56150.649 2.671 MP
GRB120805A 216.538 5.825 56144.895 3.1 MP
GRB120804A 233.948 -28.782 56143.038 1.05 MP
GRB120802A 44.843 13.768 56141.334 3.796 MP
GRB120729A 13.074 49.940 56137.456 0.8 MP
GRB120724A 245.181 3.508 56132.277 1.48 MP
GRB120722A 230.497 13.251 56130.537 0.9586 MP
GRB120716A 313.051 9.599 56124.712 2.486 MP
GRB120714B 355.409 -46.184 56122.888 0.3984 MP
GRB120712A 169.588 -20.034 56120.571 4.1745 MP
GRB120711A 94.678 -70.999 56119.114 1.405 MP
GRB120630A 352.296 42.556 56108.971 0.6 MP
GRB120624B 170.885 8.929 56102.930 2.1974 MP
GRB120521C 214.286 42.145 56068.974 6 MP
GRB120422A 136.910 14.019 56039.300 0.283 MP
GRB120404A 235.010 12.885 56021.535 2.876 MP
GRB120401A 58.082 -17.636 56018.225 4.5 [73]
GRB120327A 246.864 -29.415 56013.122 2.813 MP
GRB120326A 273.905 69.260 56012.056 1.798 MP
GRB120311A 273.092 14.296 55997.232 0.35 [64]
GRB120305A 47.536 28.492 55991.818 0.225 MP
GRB120224A 40.942 -17.761 55981.194 1.1 MP
GRB120211A 87.754 -24.775 55968.499 2.4 MP
GRB120119A 120.029 -9.082 55945.170 1.728 MP
GRB120118B 124.871 -7.185 55944.709 2.943 MP
GRB111229A 76.287 -84.711 55924.943 1.3805 MP
GRB111228A 150.067 18.298 55923.656 0.714 MP
GRB111225A 13.155 51.572 55920.160 0.297 MP
GRB111215A 349.556 32.494 55910.586 2.06 MP
GRB111211A 153.090 11.208 55906.929 0.478 MP
GRB111209A 14.344 -46.801 55904.300 0.677 MP
GRB111129A 307.434 -52.713 55894.679 1.0796 MP
GRB111123A 154.846 -20.645 55888.759 3.1516 MP
GRB111117A 12.693 23.011 55882.510 2.211 MP
GRB111107A 129.478 -66.520 55872.035 2.893 MP
GRB111008A 60.451 -32.709 55842.926 4.9898 MP
GRB111005A 223.282 -19.737 55839.337 0.0131 MP
GRB110918A 32.539 -27.105 55822.894 0.982 MP
GRB110818A 317.337 -63.981 55791.860 3.36 MP
GRB110808A 57.268 -44.194 55781.263 1.348 MP
GRB110801A 89.437 80.956 55774.826 1.858 MP
GRB110731A 280.504 -28.537 55773.465 2.83 MP
GRB110721A 333.659 -38.593 55763.200 0.382 [74, 64]
GRB110715A 237.684 -46.235 55757.551 0.82 MP
GRB110709B 164.654 -23.455 55751.898 2.109 MP
GRB110503A 132.776 52.208 55684.733 1.613 MP
GRB110422A 112.046 75.107 55673.654 1.77 MP
GRB110402A 197.402 61.253 55653.009 0.854 MP
GRB110213B 41.756 1.146 55605.605 1.083 MP
GRB110213A 42.964 49.273 55605.220 1.46 MP
GRB110205A 164.630 67.525 55597.085 2.22 MP
GRB110128A 193.896 28.065 55589.073 2.339 MP
GRB110106B 134.154 47.003 55567.893 0.618 MP
GRB110106A 79.306 64.174 55567.643 0.093 [75]
GRB101225A 0.198 44.600 55555.776 0.847 MP
GRB101224A 285.924 45.714 55554.227 0.4536 MP
GRB101219B 12.231 -34.566 55549.686 0.5519 MP
GRB101219A 74.585 -2.540 55549.105 0.718 MP
GRB101213A 241.314 21.897 55543.451 0.414 MP
GRB100906A 28.684 55.630 55445.576 1.727 MP
GRB100905A 31.550 14.930 55444.631 7.9 MP
GRB100902A 48.629 30.979 55441.814 4.5 [76]
GRB100901A 27.264 22.759 55440.565 1.408 MP
GRB100816A 351.740 26.578 55424.026 0.8035 MP
GRB100814A 22.473 -17.995 55422.160 1.44 MP
GRB100805A 299.876 52.628 55413.186 1.85 [77]
GRB100728B 44.056 0.281 55405.439 2.106 MP
GRB100728A 88.758 -15.256 55405.095 1.567 MP
GRB100724A 194.543 -11.103 55401.029 1.288 MP
GRB100704A 133.642 -24.203 55381.149 3.6 [78]
GRB100628A 225.973 -31.664 55375.345 0.102 [79, 80]
GRB100625A 15.796 -39.088 55372.773 0.452 MP
GRB100621A 315.305 -51.106 55368.127 0.542 MP
GRB100615A 177.205 -19.481 55362.083 1.398 MP
GRB100606A 350.627 -66.241 55353.800 1.5545 [81]
GRB100518A 304.789 -24.554 55334.482 4 MP
GRB100513A 169.612 3.628 55329.088 4.772 MP
GRB100508A 76.246 -20.711 55324.389 0.5201 MP
GRB100425A 299.196 -26.431 55311.119 1.755 MP
GRB100424A 209.448 1.538 55310.689 2.465 MP
GRB100418A 256.362 11.462 55304.882 0.6235 MP
GRB100414A 192.112 8.693 55300.097 1.368 MP
GRB100413A 266.221 15.834 55299.732 3.9 [82]
GRB100316D 107.628 -56.256 55271.531 0.059 MP
GRB100316B 163.488 -45.473 55271.334 1.18 MP
GRB100316A 251.978 71.827 55271.099 3.155 MP
GRB100302A 195.516 74.590 55257.829 4.813 MP
GRB100219A 154.202 -12.566 55246.636 4.6667 MP
GRB100216A 154.251 35.522 55243.422 0.038 [83, 84]
GRB100213B 124.282 43.448 55240.957 0.604 [85, 86]
GRB100206A 47.163 13.157 55233.563 0.4068 MP
GRB100117A 11.269 -1.595 55213.879 0.92 MP
GRB091208B 29.392 16.890 55173.410 1.063 MP
GRB091127 36.583 -18.952 55162.976 0.49 MP
GRB091109 309.258 -44.158 55144.207 3.076 MP
GRB091029 60.178 -55.956 55133.162 2.752 MP
GRB091024 339.248 56.890 55128.372 1.092 MP
GRB091020 175.730 50.978 55124.900 1.71 MP
GRB091018 32.186 -57.548 55122.867 0.971 MP
GRB091003 251.520 36.625 55107.191 0.8969 MP
GRB090927 343.972 -70.980 55101.422 1.37 MP
GRB090926B 46.308 -39.006 55100.914 1.24 MP
GRB090926 353.400 -66.324 55100.181 2.1062 MP
GRB090902B 264.939 27.324 55076.462 1.822 MP
GRB090814 239.610 25.631 55057.036 0.696 MP
GRB090812 353.202 -10.605 55055.251 2.452 MP
GRB090809 328.680 -0.084 55052.730 2.737 MP
GRB090726 248.680 72.884 55038.946 2.71 MP
GRB090715B 251.340 44.839 55027.877 3 MP
GRB090709 289.927 60.728 55021.319 1.8 [87]
GRB090618 293.994 78.357 55000.353 0.54 MP
GRB090530 179.419 26.594 54981.138 1.266 MP
GRB090529 212.469 24.459 54980.592 2.625 MP
GRB090519 142.279 0.180 54970.881 3.85 MP
GRB090516 138.260 -11.854 54967.352 4.109 MP
GRB090515 164.152 14.441 54966.198 0.403 [88, 89]
GRB090510 333.552 -26.583 54961.016 0.903 MP
GRB090429B 210.667 32.170 54950.229 9.4 MP
GRB090426 189.075 32.986 54947.534 2.609 MP
GRB090424 189.521 16.838 54945.589 0.544 MP
GRB090423 148.889 18.149 54944.330 8.26 MP
GRB090418 269.313 33.406 54939.464 1.608 MP
GRB090417B 209.694 47.018 54938.639 0.345 MP
GRB090407 68.980 -12.679 54928.436 1.4485 MP
GRB090404 239.240 35.516 54925.664 3 MP
GRB090401B 95.088 -8.972 54922.358 3.1 [77]
GRB090328 90.665 -41.882 54918.401 0.736 MP
GRB090323 190.710 17.053 54913.002 3.57 MP
GRB090313 198.401 8.097 54903.379 3.375 MP
GRB090205 220.911 -27.853 54867.961 4.6497 MP
GRB090201 92.052 -46.590 54863.741 2.1 MP
GRB090113 32.057 33.428 54844.778 1.7493 MP
GRB090102 128.244 33.114 54833.122 1.547 MP
GRB081230 37.331 -25.148 54830.858 2 MP
GRB081228 39.462 30.853 54828.054 3.44 MP
GRB081222 22.740 -34.095 54822.204 2.77 MP
GRB081221 15.793 -24.548 54821.681 2.26 MP
GRB081211B 168.265 53.830 54811.260 0.216 [90]
GRB081210 70.484 -11.257 54810.847 2.0631 MP
GRB081203 233.032 63.521 54803.577 2.05 MP
GRB081121 89.276 -60.603 54791.858 2.512 MP
GRB081118 82.592 -43.301 54788.623 2.58 MP
GRB081109 330.790 -54.711 54779.293 0.9787 MP
GRB081029 346.772 -68.156 54768.072 3.8479 MP
GRB081028 121.895 2.308 54767.017 3.038 MP
GRB081008 279.958 -57.431 54747.832 1.9685 MP
GRB081007 339.960 -40.147 54746.224 0.5295 MP
GRB080928 95.070 -55.200 54737.626 1.692 MP
GRB080916C 119.847 -56.638 54725.009 4.35 MP
GRB080916 336.276 -57.023 54725.406 0.689 MP
GRB080913 65.728 -25.130 54722.283 6.695 MP
GRB080906 228.044 -80.518 54715.565 2.1 MP
GRB080905B 301.741 -62.563 54714.705 2.374 MP
GRB080905 287.674 -18.880 54714.499 0.1218 MP
GRB080825B 209.201 -68.955 54703.741 4.3 MP
GRB080810 356.793 0.320 54688.549 3.35 MP
GRB080805 314.223 -62.445 54683.321 1.505 MP
GRB080804 328.668 -53.185 54682.972 2.2045 MP
GRB080721 224.483 -11.724 54668.434 2.591 MP
GRB080710 8.274 19.502 54657.301 0.845 MP
GRB080707 32.618 33.109 54654.353 1.23 MP
GRB080607 194.947 15.920 54624.255 3.036 MP
GRB080605 262.125 4.016 54622.992 1.6398 MP
GRB080604 236.965 20.558 54621.310 1.416 MP
GRB080603B 176.532 68.061 54620.818 2.69 MP
GRB080603 279.409 62.744 54620.471 1.688 MP
GRB080602 19.176 -9.232 54619.063 1.8204 MP
GRB080520 280.193 -54.992 54606.931 1.545 MP
GRB080517 102.242 50.735 54603.891 0.089 MP
GRB080516 120.642 -26.160 54602.012 3.2 [91]
GRB080515 3.163 32.578 54601.251 2.47 MP
GRB080514B 322.845 0.708 54600.414 1.8 MP
GRB080430 165.311 51.686 54586.828 0.767 MP
GRB080413B 326.144 -19.981 54569.369 1.1 MP
GRB080413 287.299 -27.678 54569.121 2.433 MP
GRB080411 37.980 -71.302 54567.886 1.03 MP
GRB080330 169.269 30.623 54555.154 1.51 MP
GRB080325 277.893 36.524 54550.173 1.78 MP
GRB080319C 258.981 55.392 54544.518 1.95 MP
GRB080319B 217.921 36.302 54544.259 0.937 MP
GRB080319 206.333 44.080 54544.240 2.0265 MP
GRB080310 220.058 -0.175 54535.360 2.42 MP
GRB080210 251.267 13.827 54506.326 2.641 MP
GRB080207 207.512 7.502 54503.896 2.0858 MP
GRB080205 98.253 62.792 54501.330 2.72 MP
GRB080129 105.284 -7.846 54494.255 4.349 MP
GRB080123 338.943 -64.901 54488.182 0.495 MP
GRB071227 58.130 -55.984 54461.843 0.383 MP
GRB071122 276.605 47.075 54426.058 1.14 MP
GRB071117 335.044 -63.443 54421.618 1.331 MP
GRB071112C 39.212 28.371 54416.773 0.823 MP
GRB071031 6.405 -58.060 54404.046 2.692 MP
GRB071028B 354.163 -31.621 54401.114 0.94 [92]
GRB071025 355.071 31.778 54398.173 5.2 MP
GRB071021 340.643 23.718 54394.404 2.452 MP
GRB071020 119.665 32.861 54393.293 2.145 MP
GRB071010B 150.539 45.730 54383.865 0.947 MP
GRB071010 288.060 -32.402 54383.154 0.98 MP
GRB071003 301.851 10.947 54376.320 1.6043 MP
GRB070810 189.963 10.751 54322.092 2.17 MP
GRB070809 203.770 -22.142 54321.807 0.2187 MP
GRB070802 36.899 -55.528 54314.297 2.45 MP
GRB070724 27.808 -18.594 54305.454 0.457 MP
GRB070721B 33.137 -2.195 54302.440 3.626 MP
GRB070714B 57.843 28.298 54295.208 0.92 MP
GRB070714 42.930 30.243 54295.139 1.58 MP
GRB070612 121.369 37.269 54263.110 0.617 MP
GRB070611 1.992 -29.756 54262.081 2.04 MP
GRB070529 283.742 20.659 54249.534 2.4996 MP
GRB070521 242.661 30.256 54241.286 2.0865 MP
GRB070518 254.199 55.295 54238.602 1.161 MP
GRB070508 312.799 -78.385 54228.179 0.82 [93, 94]
GRB070506 347.218 10.722 54226.233 2.31 MP
GRB070429B 328.015 -38.829 54219.131 0.904 MP
GRB070419B 315.707 -31.264 54209.447 1.9591 MP
GRB070419 182.745 39.925 54209.416 0.97 MP
GRB070411 107.333 1.064 54201.842 2.954 MP
GRB070328 65.115 -34.067 54187.162 2.0627 MP
GRB070318 48.487 -42.946 54177.312 0.836 MP
GRB070306 148.097 10.482 54165.695 1.4965 MP
GRB070224 179.027 -13.330 54155.853 1.9922 MP
GRB070223 153.452 43.134 54154.052 1.6295 MP
GRB070208 197.886 61.965 54139.382 1.165 MP
GRB070129 37.004 11.684 54129.983 2.3384 MP
GRB070125 117.824 31.151 54125.306 1.547 MP
GRB070110 0.913 -52.974 54110.307 2.352 MP
GRB070103 352.558 26.876 54103.866 2.6208 MP
GRB061222B 105.353 -25.860 54091.174 3.355 MP
GRB061222 358.264 46.533 54091.145 2.088 MP
GRB061217 160.413 -21.124 54086.153 0.827 MP
GRB061210 144.522 15.621 54079.514 0.4097 MP
GRB061202 105.525 -74.698 54071.341 2.2543 MP
GRB061201 332.134 -74.580 54070.666 0.111 MP
GRB061126 86.602 64.211 54065.367 1.1588 MP
GRB061121 147.227 -13.195 54060.641 1.314 MP
GRB061110B 323.919 6.876 54049.916 3.44 MP
GRB061110 336.291 -2.258 54049.491 0.758 MP
GRB061021 145.150 -21.952 54029.652 0.3463 MP
GRB061007 46.332 -50.501 54015.422 1.261 MP
GRB061006 111.031 -79.199 54014.699 0.4377 MP
GRB061002 220.347 48.742 54010.044 0.564 [67]
GRB060927 329.550 5.364 54005.589 5.47 MP
GRB060926 263.932 13.038 54004.700 3.204 MP
GRB060923B 238.195 -30.904 54001.485 1.5094 MP
GRB060923 254.617 12.361 54001.217 2.6 [95, 87]
GRB060912 5.284 20.972 53990.580 0.937 MP
GRB060908 31.826 0.342 53986.373 1.8836 MP
GRB060906 40.754 30.362 53984.356 3.6855 MP
GRB060904B 58.211 -0.725 53982.105 0.703 MP
GRB060814 221.339 20.586 53961.960 1.9229 MP
GRB060805 220.931 12.586 53952.200 2.3633 MP
GRB060801 213.006 16.982 53948.511 1.131 MP
GRB060729 95.382 -62.370 53945.800 0.54 MP
GRB060719 18.432 -48.381 53935.285 1.532 MP
GRB060714 227.860 -6.566 53930.633 2.7105 MP
GRB060708 7.808 -33.759 53924.511 1.92 MP
GRB060707 357.080 -17.905 53923.896 3.425 MP
GRB060614 320.884 -53.027 53900.530 0.125 MP
GRB060607 329.710 -22.496 53893.217 3.0749 MP
GRB060605 322.156 -6.059 53891.761 3.773 MP
GRB060604 337.229 -10.916 53890.763 2.1357 MP
GRB060602 149.570 0.304 53888.897 0.787 MP
GRB060526 232.826 0.285 53881.686 3.221 MP
GRB060522 322.937 2.886 53877.091 5.11 MP
GRB060512 195.774 41.191 53867.968 2.1 MP
GRB060510B 239.122 78.570 53865.349 4.94 MP
GRB060510 95.867 -1.163 53865.322 1.2 [77]
GRB060505 331.764 -27.815 53860.275 0.089 MP
GRB060502B 278.938 52.632 53857.725 0.287 MP
GRB060502 240.927 66.601 53857.127 1.51 MP
GRB060418 236.428 -3.639 53843.129 1.4895 MP
GRB060319 176.388 60.011 53813.039 1.172 MP
GRB060306 41.095 -2.148 53800.034 1.559 MP
GRB060223 55.206 -17.130 53789.253 4.41 MP
GRB060218 50.416 16.867 53784.149 0.0331 MP
GRB060210 57.739 27.026 53776.208 3.91 MP
GRB060206 202.931 35.051 53772.199 4.048 MP
GRB060204B 211.812 27.677 53770.607 2.3393 MP
GRB060202 35.846 38.384 53768.362 0.783 MP
GRB060124 77.108 69.741 53759.663 2.296 MP
GRB060123 179.700 45.514 53758.932 0.56 MP
GRB060121 137.466 45.663 53756.933 4.6 [96]
GRB060116 84.693 -5.437 53751.359 6.6 [97]
GRB060115 54.035 17.345 53750.547 3.53 MP
GRB060111 276.205 37.604 53746.183 2.32 MP
GRB051227 125.241 31.925 53731.755 0.8 MP
GRB051221 328.703 16.891 53725.077 0.5468 MP
GRB051210 330.171 -57.614 53714.241 2.5 MP
GRB051117B 85.181 -19.274 53691.558 0.481 MP
GRB051111 348.138 18.375 53685.250 1.5495 MP
GRB051109B 345.460 38.680 53683.361 0.08 MP
GRB051109 330.314 40.823 53683.050 2.346 MP
GRB051022 359.017 19.607 53665.547 0.809 MP
GRB051016B 132.116 13.656 53659.770 0.9364 MP
GRB051008 202.873 42.098 53651.690 2.77 MP
GRB051006 110.809 9.506 53649.855 1.059 MP
GRB051001 350.953 -31.523 53644.466 2.4296 MP
GRB050922C 317.388 -8.758 53635.830 2.198 MP
GRB050922B 5.806 -5.605 53635.626 4.5 MP
GRB050915 81.687 -28.016 53628.474 2.5273 MP
GRB050911 13.657 -38.849 53624.666 0.165 [98, 99]
GRB050908 20.461 -12.955 53621.238 3.344 MP
GRB050904 13.711 14.085 53617.078 6.29 MP
GRB050826 87.756 -2.643 53608.263 0.297 MP
GRB050824 12.234 22.609 53606.967 0.83 MP
GRB050822 51.113 -46.033 53604.159 1.434 MP
GRB050820 337.409 19.560 53602.274 2.612 MP
GRB050819 358.757 24.861 53601.683 2.5043 MP
GRB050814 264.189 46.339 53596.485 5.77 MP
GRB050813 241.989 11.248 53595.281 0.72 [100, 101]
GRB050803 350.658 5.786 53585.801 3.5 MP
GRB050802 219.274 27.787 53584.422 1.71 MP
GRB050801 204.146 -21.928 53583.769 1.38 MP
GRB050730 212.072 -3.772 53581.832 3.9678 MP
GRB050724 246.185 -27.541 53575.524 0.2575 MP
GRB050714B 169.699 -15.547 53565.945 2.4383 MP
GRB050709 345.362 -38.978 53560.942 0.16 MP
GRB050525 278.136 26.339 53515.002 0.606 MP
GRB050509B 189.058 28.984 53499.167 0.2255 MP
GRB050505 141.764 30.273 53495.974 4.27 MP
GRB050502B 142.542 16.997 53492.393 5.2 MP
GRB050502 202.443 42.674 53492.093 3.793 MP
GRB050416 188.478 21.057 53476.462 0.6535 MP
GRB050408 180.573 10.852 53468.683 1.2357 MP
GRB050406 34.468 -50.188 53466.666 2.44 [102]
GRB050401 247.870 2.187 53461.597 2.9 MP
GRB050319 154.200 43.548 53448.397 3.24 MP
GRB050318 49.713 -46.396 53447.656 1.44 MP
GRB050315 306.476 -42.600 53444.875 1.949 MP
GRB050223 271.385 -62.472 53424.131 0.5915 MP
GRB050219 166.412 -40.683 53420.528 0.211 MP
GRB050215B 174.448 40.797 53416.107 2.62 MP
GRB050126 278.113 42.370 53396.501 1.29 MP
GRB041006 13.709 1.235 53284.512 0.716 MP
GRB040924 31.594 16.114 53272.494 0.859 MP
GRB040912 359.226 -1.001 53260.592 1.563 MP
GRB040701 311.943 -40.237 53189.583 0.2146 [103, 104]
GRB031203 120.626 -39.850 52976.918 0.105 MP
GRB030528 256.008 -22.650 52787.544 0.782 MP
GRB030429 183.281 -20.914 52758.446 2.656 MP
GRB030329 161.208 21.522 52727.484 0.1685 MP
GRB030328 182.702 -9.348 52726.473 1.52 MP
GRB030323 166.539 -21.770 52721.915 3.372 MP
GRB030226 173.271 25.898 52696.157 1.986 MP
GRB030115 169.625 15.001 52654.141 2.5 [105]
GRB021211 122.270 6.678 52619.471 1.004 MP
GRB021004 6.745 18.949 52551.504 2.33 MP
GRB020903 342.176 -20.769 52520.421 0.25 MP
GRB020819B 351.840 6.250 52505.623 1.9621 MP
GRB020813 296.658 -19.588 52499.114 1.2545 MP
GRB020405 209.513 -31.373 52369.029 0.69 MP
GRB020305 190.616 -14.303 52338.629 0.2 MP
GRB020127 123.756 36.776 52301.873 1.9 MP
GRB020124 143.211 -11.520 52298.445 3.198 MP
GRB011211 168.825 -21.949 52254.798 2.14 MP
GRB011121 173.606 -76.028 52234.783 0.36 MP
GRB010921 344.000 40.931 52173.219 0.45 MP
GRB010222 223.052 43.018 51962.308 1.477 MP
GRB000926 256.040 51.786 51813.877 2.066 MP
GRB000911 34.681 7.741 51798.302 1.0585 MP
GRB000607 38.496 17.002 51702.101 0.1405 [106]
GRB000418 186.330 20.103 51652.896 1.118 MP
GRB000301C 245.078 29.443 51604.529 2.03 MP
GRB000214 283.613 -66.458 51588.042 0.42 MP
GRB000210 29.815 -40.659 51584.364 0.8463 MP
GRB000131 93.379 -51.944 51574.624 4.5 MP
GRB991216 77.380 11.285 51528.672 1.02 MP
GRB991208 248.473 46.456 51520.192 0.706 MP
GRB990712 337.971 -73.408 51371.323 0.434 MP
GRB990705 $7$72.131 51364.668 0.842 MP
GRB990510 204.525 -80.500 51308.367 1.619 MP
GRB990506 178.709 -26.676 51304.475 1.3 MP
GRB990123 231.374 44.758 51201.408 1.6 MP
GRB981226 352.405 -23.932 51173.408 1.11 [107]
GRB980703 359.779 8.585 50997.182 0.966 MP
GRB980613 154.442 71.486 50977.202 1.096 MP
GRB980425 293.725 -52.819 50928.446 0.0085 MP
GRB980329 105.658 38.846 50901.156 3.5 MP
GRB980326 129.143 -18.857 50898.888 1 [108]
GRB971214 179.110 65.200 50796.110 3.42 MP
GRB970828 272.106 59.302 50688.739 0.9578 MP
GRB970508 103.456 79.272 50576.904 0.835 MP
GRB970228 75.488 11.768 50507.124 0.695 MP
Appendix B GRBs relevant for Figure 1
In this appendix we describe some properties of the GRB-neutrino events appearing in Fig. 1.
As already stressed in the main text, our analysis relied on the latest HESE neutrino data release by IceCube [39]. We considered only shower neutrinos, for which the energy is approximately equal to the visible energy [41, 42] (whereas the energy of track neutrinos is poorly known).
The GRBs which we analyzed (also in producing simulated data) are those of our catalogue reported in Appendix A. It is worth emphasizing that our catalogue agrees with both [44] and [45] for what concerns the four GRBs relevant for Fig. 1.
[…]